# On the asymptotics of orthogonal polynomials on multiple intervals with non-analytic weights

### v2

@@ 2.3. Construction of the global parametrix
$\tilde{\mathbf{T}}_{n}^{+}(z)$ $=\tilde{\mathbf{T}}_{n}^{-}(z)\begin{cases}\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2n\mathfrak{g}(z)}&1\end{bmatrix}&z\in\Sigma_{j},\vspace{3pt}\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in(a_{j},b_{j}),\vspace{3pt}\\ \,\mathrm{e}^{-\Delta_{j}{\boldsymbol{\sigma}_{3}}}&z\in(b_{j},a_{j+1}),\end{cases}$
@@
$\mathbf{U}^{+}_{n}(z;\mu)=\begin{cases}\mathbf{U}^{-}_{n}(z;\mu)\begin{bmatrix}1&0\\ f_{j}^{\sharp}(x,y)/\omega_{j}(z)\,\mathrm{e}^{-2(n\mathfrak{g}(z)-G(z))}&1\end{bmatrix}&z\in\Sigma_{j}\setminus\mathbb{R},\vspace{4pt}\\ \mathbf{U}^{-}_{n}(z;\mu)\begin{bmatrix}0&1\\ -1&0\end{bmatrix}&z\in(a_{j},b_{j}),\vspace{4pt}\\ \mathbf{U}^{-}_{n}(z;\mu))\begin{bmatrix}\,\mathrm{e}^{-n\Delta_{j}-\zeta_{j}}&0\\ 0&\,\mathrm{e}^{n\Delta_{j}+\zeta_{j}}\end{bmatrix}&z\in(b_{j},a_{j+1}),\end{cases}$
@@
For a given $z\in\mathbb{C}\setminus[a,b]$ let $\eta(z)$ be the closet point in $[a,b]$ to $z$. Set $S(z)=\sqrt{z-a}\sqrt{z-b}$. Then
@@ 2.4. Using local solutions
From the jump conditions established in Appendix C, we find for $|z-b_{j}|<\epsilon$, the function
$\mathbf{R}_{n}(z)=\mathbf{T}_{n}(z)(\tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}\mathbf{P}_{n}(z;b_{j}))^{-1},$
is analytic for $z\not\in(b_{j}-\epsilon,b_{j})$, and for $|z-b_{j}|<\delta$. Then we compute, for $b_{j}-\delta<z<b_{j}$
@@
×$\times\left(\frac{\tilde{\rho}(z;\delta)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{Q}_{n}^{-}(z;b_{j})^{-1}\Psi(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\tilde{\rho}(z;\delta)}{\rho(z;b_{j})}\right)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi(z)^{\boldsymbol{\sigma}_{3}}\mathbf{A}^{-}_{n}(z;b_{j})^{-1}\mathbf{G}_{n}^{-}(z)\tilde{\mathbf{G}}_{n}^{-}(z)^{-1}.$
@@
$\mathbf{B}_{2}(z)$ $=\mathbf{Q}_{n}^{-}(z;b_{j})^{-1}\Psi(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E},$
$\mathbf{B}_{3}(z)$ $=\mathbf{E}^{-1}\Psi(z)^{\boldsymbol{\sigma}_{3}}\mathbf{A}^{-}_{n}(z;b_{j})^{-1}\mathbf{G}_{n}^{-}(z)\tilde{\mathbf{G}}_{n}^{-}(z)^{-1}.$
@@ 2.5. Global approximation
$\mathbf{R}_{n}(x,y)=\mathbf{T}_{n}(z,y)\begin{cases}(\tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}\mathbf{P}_{n}(z;b_{j}))^{-1},&|z-b_{j}|<\epsilon,\\ \mathbf{(}\tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}\mathbf{P}_{n}(z;a_{j}))^{-1},&|z-a_{j}|<\epsilon,\\ \mathbf{G}_{n}(z)^{-1}&\text{otherwise}.\end{cases}$
@@
¯$\overline{\partial}\mathbf{R}_{n}(x,y)=\mathbf{R}_{n}(x,y)\mathbf{X}_{n}(x,y),$
$\mathbf{X}_{n}(x,y)=\begin{cases}\mathbf{G}_{n}(z)\mathbf{W}_{n}(x,y)\mathbf{G}_{n}(z)^{-1}&z\in\Omega_{j,\pm}\cap(\overline{B_{\epsilon}(a_{j})\cup B_{\epsilon}(b_{j})})^{c},\vspace{4pt}\\ \tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}\mathbf{P}_{n}(z;c)\mathbf{W}_{n}(x,y)\mathbf{P}_{n}(z;c)^{-1}\mathbf{G}_{n}(z)\tilde{\mathbf{G}}_{n}(z)^{-1}&z\in\Omega_{j,\pm}\cap B_{\epsilon}(c),\quad c=a_{j},b_{j},\vspace{4pt}\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}),\end{cases}$
@@
The jump condition on $(b_{j}-\epsilon,b_{j}-\delta)$, for example, comes from Lemma 2.2, and the jump condition on $|z-c|=\epsilon$ follows from (2.6) and the fact that $\mathbf{G}_{n}^{-1}\mathbf{P}_{n}(z;c)=\mathbf{I}+\mathrm{O}(n^{-1})$ for $|z-c|=\epsilon$, see Appendix C.
@@
¯$\overline{\partial}\mathbf{C}_{n}(x,y)$ $=\mathbf{C}_{n}(z)\mathbf{X}_{n}(x,y),\quad z\in\Omega_{j,+}\cup\Omega_{j,-}$
@@
$2p(|\alpha_{j}|-k)+p-4<0,\quad 2p(|\beta_{j}|-k)+p-4<0,$
@@
$\|\mathbf{X}_{n}\|_{L^{p}(\Omega)}=\mathrm{O}\left(n^{2(|\alpha_{j}|-k)+1-\frac{4}{p}}+n^{-\frac{(3+k)p-1)((2k-1)p+4)}{(2k+5)p^{2}}}+\delta^{k}\right).$
@@
where $c$ is sufficiently small, but fixed, such that the estimates in Lemma C.1 hold, and $N\ll n^{2}$. In $\Omega_{b_{j},\mathrm{I}}$ and $\Omega_{b_{j},\mathrm{II}}$ we simply estimate $|\,\mathrm{e}^{-\mathfrak{g}(z)}|\leq 1$. Then for $\alpha_{j}\neq 0$, $z\in\Omega_{j,\mathrm{I}}$ we have for $D_{\mathrm{I}}>0$, using Lemma C.1,
@@
$\int_{\Omega_{b_{j},\mathrm{I}}}\|\mathbf{X}_{n}(x,y)\|^{p}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y\leq D_{\mathrm{I}}^{p}n^{2p(|\alpha_{j}|-k)+p-4}\int_{0}^{2\pi}\int_{0}^{c}r^{-p|\alpha_{j}|+pk+1}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\theta+O(n^{-2}\delta^{kp}).$
@@
$p(k-|\alpha_{j})+1>-1,\quad 2p(|\alpha_{j}|-k)+p-4<0,$
and the second is more restrictive. Now, if $\alpha_{j}=0$ we have
@@
$\int_{\Omega_{b_{j},\mathrm{I}}}\|\mathbf{X}_{n}(x,y)\|^{p}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y$ $\leq D_{\mathrm{I}}^{p}n^{-2pk+p-4}(\log n)^{2}+O(n^{-2}\delta^{kp}),$
@@
Then, again, rescaling and using polar coordinates
$\int_{\Omega_{b_{j},\mathrm{II}}}\|\mathbf{X}_{n}(x,y)\|^{p}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y$ $\leq D_{\mathrm{II}}^{p}\left[N^{3p/2}n^{-4-2kp}\int_{0}^{N}r^{kp+1}\mathop{}\!\mathrm{d}r\right]+\mathrm{O}(N/n^{2}\delta^{kp}),$
$=\frac{D_{\mathrm{II}}^{p}}{kp+2}N^{3p/2+kp+2}n^{-4-2kp}+\mathrm{O}(N/n^{2}\delta^{kp}).$
@@
Suppose $p>2$ and $p(k-|\alpha_{j}|)+2>0,p(k-|\beta_{j}|)+2>0$ for all $j$. Suppose also that $\mathbf{V}_{n}(x,y)\in L^{\infty}(\mathbb{C})$ satisfies
@@
Then $\mathbf{V}_{n}(x,y)+\mathbf{I}$ is a solution of the $\overline{\partial}$ problem solved by $\mathbf{C}_{n}(x,y).$ Lastly, $\mathbf{V}_{n}$ is $\alpha$-Hölder continuous for some $10<\alpha<$ 1 and $\mathbf{V}_{n}|_{\Omega}$ is differentiable almost everywhere.
@@
We use [1, Theorem 4.3.10] which states that, in a distributional sense, the integral operator
$\mathcal{K}u(x,y)=\frac{1}{\pi}\int_{\Omega}\frac{u(x^{\prime},y^{\prime})}{z^{\prime}-z}\mathop{}\!\mathrm{d}A(z^{\prime}),$
@@
¯$\overline{\partial}\mathbf{V}_{n}$ $=\mathbf{V}_{n}\mathbf{X}_{n}+\mathbf{X}_{n},$
¯$\overline{\partial}(\mathbf{V}_{n}+\mathbf{I})$ $=(\mathbf{V}_{n}+\mathbf{I})\mathbf{X}_{n}.$
Next, from [1, Theorem 4.3.13], we have that $\mathcal{K}$ maps $L^{p}(\mathbb{C})$ into the space of $\alpha$-Hölder continuous functions for $\alpha=1-2/p$, $p>2$. And we again use the relation
$\mathbf{V}_{n}=\mathcal{K}(\mathbf{V}_{n}\mathbf{X}_{n})+\mathcal{K}\mathbf{X}_{n},$
to get the desired conclusion because $\mathbf{X}_{n}\in L^{p}(\Omega)$. The claim about $\mathbf{V}_{n}|_{\Omega}$ follows from the elliptic regularity theorem, see [20, Theorem 9.26], since $\mathbf{V}_{n}|_{\Omega}\in L^{\infty}(\Omega)$. ∎
We also see that for any $p>2>q,1/p+1/q=1$,
@@
Fix $k\geq 1$ and set $\eta=\max_{j}\max\{|\alpha_{j}|,|\beta_{j}|\}$. Suppose $02(\eta-k)-1<$ 0. Then for every $\gamma>0$, and $n$ sufficiently large, there is a unique $L^{\infty}(\mathbb{C})$ solution of (2.8) that satisfies
@@
If $p=2$ is Lemma 2.3, the bound is $\mathrm{O}(n^{-2k-1}+n^{2(\eta-k)-1}+\delta^{k})$ and by choosing $p>2$, sufficiently close to $2$, the theorem follows by Proposition 2.4. ∎
@@
For $k\geq 1$ such that $12(\eta-k)<$ 1, $\eta=\max_{j}\max\{|\alpha_{j}|,|\beta_{j}|\}$ and any fixed $\gamma>0$, sufficiently small, set
$\mathbf{D}^{\natural}_{j}(z;c)$ $=\begin{bmatrix}d_{j,1}^{\sharp}(x,y;c)+1&0\\ 0&d_{j,-1}^{\sharp}(x,y;c)+1\end{bmatrix},$
$\mathbf{D}^{\natural}(x,y)$ =$=\begin{cases}\mathbf{D}^{\natural}_{j}(z;a_{j})^{-1}\mathbf{D}^{\natural}_{j}(z;b_{j})^{-1},&z\in\Omega_{j,\pm},\\ \mathbf{I}&\text{otherwise},\end{cases}$
$\mathbf{F}^{\natural}(x,y)$ =$=\begin{cases}\begin{bmatrix}1&0\\ \pm f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,\pm},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
@@
All the estimates have been previously discussed and the remaining issue is the equality — that $\tilde{\mathbf{Y}}_{n}(z;\mu)=\mathbf{Y}_{n}(z;\mu)$. It can be shown that the jump conditions for $\tilde{\mathbf{R}}_{n}(z)$ satisfy the so-called product condition [33, Definition 2.55] and derivatives of the jump matrix are essentially bounded. Thus $\mathbf{O}_{n}\in H^{1}_{z}(\Gamma_{\delta})$, [33, Definition 2.48], an appropriate Sobolev space to ensure that $\tilde{\mathbf{R}}_{n}(z)$, as defined, is uniformly $\alpha$-Hölder continuous and takes $\alpha$-Hölder continuous boundary values for some $\alpha>0$. And $\tilde{\mathbf{Y}}_{n}(z;\mu)$ satisfies the jump conditions set for the FIK RH problem, in a continuous sense (away from $a_{j}$, $b_{j}$).
Next, we need to verify analyticity. But from classical elliptic regularity, [20, Theorem 9.26], we have that $\tilde{\mathbf{Y}}_{n}(\cdot;\mu)\in C^{\infty}(\Omega^{\prime})$ for every open set $\Omega^{\prime}$ that does not intersect any one of the contours used in the deformation, i.e. for
@@
Furthermore, ¯$\overline{\partial}\tilde{\mathbf{Y}}_{n}=0$ in such a set $\Omega^{\prime}$, implying analyticity.
@@
So, if $n^{-1}\delta^{-3/2}\ll 1$, or $\delta=n^{-2/3+\gamma}$, and $02(\eta-k)+1/3<$ 0 we obtain a valid error term giving uniform asymptotics on scaled neighborhoods of the endpoints $a_{j},b_{j}$. Furthermore, by taking $k$ sufficiently large, $\gamma<2/3$ here can be chosen arbitrarily close to $2/3$.

### v4

@@ 2.3. Construction of the global parametrix
$\tilde{\mathbf{T}}_{n}^{+}(z)$ $=\tilde{\mathbf{T}}_{n}^{-}(z)\begin{cases}\begin{bmatrix}1&0\\ \,\mathrm{e}^{-2n\mathfrak{g}(z)}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j},\vskip 3.0pt\\ \begin{bmatrix}0&\rho(x)\\ -1/\rho(x)&0\end{bmatrix}&z\in(a_{j},b_{j}),\vskip 3.0pt\\ \,\mathrm{e}^{-\Delta_{j}{\boldsymbol{\sigma}_{3}}}&z\in(b_{j},a_{j+1}),\end{cases}$
@@
$\mathbf{U}^{+}_{n}(z;\mu)=\begin{cases}\mathbf{U}^{-}_{n}(z;\mu)\begin{bmatrix}1&0\\ \,\mathrm{e}^{-2(n\mathfrak{g}(z)-G(z))}f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Sigma_{j}\setminus\mathbb{R},\vskip 4.0pt\\ \mathbf{U}^{-}_{n}(z;\mu)\begin{bmatrix}0&1\\ -1&0\end{bmatrix}&z\in(a_{j},b_{j}),\vskip 4.0pt\\ \mathbf{U}^{-}_{n}(z;\mu))\begin{bmatrix}\,\mathrm{e}^{-n\Delta_{j}-\zeta_{j}}&0\\ 0&\,\mathrm{e}^{n\Delta_{j}+\zeta_{j}}\end{bmatrix}&z\in(b_{j},a_{j+1}),\end{cases}$
@@
Note that the constants $\zeta_{j}$ are the same as those in the definition of $G(z)$. This is so that the ratio $G(z)/\tilde{G}(z)$ will be analytic across the gaps $(a_{j+1},b_{j})$.
For a given $z\in\mathbb{C}\setminus[a,b]$ let $\eta(z)$ be the closest point in $[a,b]$ to $z$. Set $S(z)=\sqrt{z-a}\sqrt{z-b}$. Then
@@ 2.4. Using local solutions
In this section we make heavy use of the definitions and jump conditions established in Appendix C. Consider the function
$\mathbf{R}_{n}(x,t)=\mathbf{T}_{n}(x,y)(\tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}\mathbf{P}_{n}(z;b_{j}))^{-1}.$
We claim that for $|z-b_{j}|<\epsilon$, that the function $\mathbf{R}_{n}$ is continuous for $z\not\in(b_{j}-\epsilon,b_{j})$, and for $|z-b_{j}|<\delta$. Indeed, $\mathbf{T}_{n}(x,y)$ and $\mathbf{P}_{n}(z;b_{j})$ have the same jump condition for $z\not\in(b_{j}-\epsilon,b_{j})$ and $\tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}$ is analytic for $z\not\in(b_{j}-\epsilon,b_{j})$.
We then recall that for $b_{j}-\epsilon<z<b_{j}$ we have
${\mathbf{T}}_{n}^{+}(x,y)$ $={\mathbf{T}}_{n}^{-}(x,y)\begin{bmatrix}0&\tilde{\rho}(z;b_{j})\\ -1/\tilde{\rho}(z;b_{j})&0\end{bmatrix}$.
Using that for $b_{j}-\epsilon<z<b_{j}$
$\mathbf{A}_{n}^{+}(z;b_{j})$ $=\mathbf{A}_{n}^{-}(z;b_{j})\Psi^{+}(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\rho(z)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi^{+}(z)^{\boldsymbol{\sigma}_{3}},$
and
$\mathbf{Q}_{n}^{+}(z;b_{j})$ $=\mathbf{Q}_{n}^{-}(z;b_{j})\begin{bmatrix}0&\rho(z;b_{j})\\ -1/\rho(z;b_{j})&0\end{bmatrix},\quad\quad b_{j}-\epsilon<z<b_{j},$
we compute, again for $b_{j}-{\epsilon}<z<b_{j}$, that
@@
$\times\left(\frac{\tilde{\rho}(z;\delta)}{\rho(z;b_{j})}\right)^{{\boldsymbol{\sigma}_{3}}}\mathbf{Q}_{n}^{-}(z;b_{j})^{-1}\Psi^{+}(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}\left(\frac{\tilde{\rho}(z;\delta)}{\rho(z;b_{j})}\right)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E}^{-1}\Psi^{+}(z)^{\boldsymbol{\sigma}_{3}}\mathbf{A}^{-}_{n}(z;b_{j})^{-1}\mathbf{G}_{n}^{-}(z)\tilde{\mathbf{G}}_{n}^{-}(z)^{-1}.$
For $b_{j}-\delta<z<b_{j}$ we have that $\tilde{\rho}(z;\delta)=\rho(z;b_{j})$ and the claims about $\mathbf{R}_{n}$ follow.
@@
$\mathbf{B}_{2}(z)$ $=\mathbf{Q}_{n}^{-}(z;b_{j})^{-1}\Psi^{+}(z)^{-{\boldsymbol{\sigma}_{3}}}\mathbf{E},$
$\mathbf{B}_{3}(z)$ $=\mathbf{E}^{-1}\Psi^{+}(z)^{\boldsymbol{\sigma}_{3}}\mathbf{A}^{-}_{n}(z;b_{j})^{-1}\mathbf{G}_{n}^{-}(z)\tilde{\mathbf{G}}_{n}^{-}(z)^{-1}.$
@@ 2.5. Global approximation
$\mathbf{R}_{n}(x,y)=\mathbf{T}_{n}(x,y)\begin{cases}(\tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}\mathbf{P}_{n}(z;b_{j}))^{-1},&|z-b_{j}|<\epsilon,\\ \mathbf{(}\tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}\mathbf{P}_{n}(z;a_{j}))^{-1},&|z-a_{j}|<\epsilon,\\ \mathbf{G}_{n}(z)^{-1}&\text{otherwise}.\end{cases}$
@@
$\overline{\partial}\mathbf{R}_{n}(x,y)=\mathbf{R}_{n}(x,y)\mathbf{X}_{n}(x,y),$
$\mathbf{X}_{n}(x,y)=\begin{cases}\mathbf{G}_{n}(z)\mathbf{W}_{n}(x,y)\mathbf{G}_{n}(z)^{-1}&z\in\Omega_{j,\pm}\cap(\overline{B_{\epsilon}(a_{j})\cup B_{\epsilon}(b_{j})})^{c},\vskip 4.0pt\\ \tilde{\mathbf{G}}_{n}(z)\mathbf{G}_{n}(z)^{-1}\mathbf{P}_{n}(z;c)\mathbf{W}_{n}(x,y)\mathbf{P}_{n}(z;c)^{-1}\mathbf{G}_{n}(z)\tilde{\mathbf{G}}_{n}(z)^{-1}&z\in\Omega_{j,\pm}\cap B_{\epsilon}(c),\quad c=a_{j},b_{j},\vskip 4.0pt\\ \mathbf{0}&z\in\mathbb{C}\setminus(\overline{\Omega}_{j,+}\cup\overline{\Omega}_{j,-}),\end{cases}$
@@
The jump condition on $(b_{j}-\epsilon,b_{j}-\delta)$, for example, comes from Lemma 2.2, and the jump condition on $|z-c|=\epsilon$ follows from (2.6) and the fact that $\mathbf{G}_{n}^{-1}(z)\mathbf{P}_{n}(z;c)=\mathbf{I}+\mathrm{O}(n^{-1})$ for $|z-c|=\epsilon$, see Appendix C.
@@
$\overline{\partial}\mathbf{C}_{n}(x,y)$ $=\mathbf{C}_{n}(z)\mathbf{X}_{n}(x,y),\quad z\in\Omega_{j,+}\cup\Omega_{j,-}$
@@
$2p({2}|\alpha_{j}|-k)+p-4<0,\quad 2p({2}|\beta_{j}|-k)+p-4<0,$
@@
$\|\mathbf{X}_{n}\|_{L^{p}(\Omega)}=\mathrm{O}\left(n^{2(\zeta-k)+1-\frac{4}{p}}+n^{-\frac{((3+k)p-1)((2k-1)p+4)}{(2k+5)p^{2}}}+\delta^{k}\right),\quad\zeta=2\max_{j}\max\{|\alpha_{j}|,|\beta_{j}|\}.$
@@
where $c$ is sufficiently small, but fixed, such that the estimates in Lemma C.1 hold, and $N\ll n^{2}$. In $\Omega_{b_{j},\mathrm{I}}$ and $\Omega_{b_{j},\mathrm{II}}$ we simply estimate $|\,\mathrm{e}^{-\mathfrak{g}(z)}|\leq 1$. Then for $\alpha_{j}\neq 0$, $z\in\Omega_{{b_{j}},\mathrm{I}}$ we have for $D_{\mathrm{I}}>0$, using Lemma C.1,
@@
$\int_{\Omega_{b_{j},\mathrm{I}}}\|\mathbf{X}_{n}(x,y)\|^{p}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y\leq D_{\mathrm{I}}^{p}n^{2p({2}|\alpha_{j}|-k)+p-4}\int_{0}^{2\pi}\int_{0}^{c}r^{-p|\alpha_{j}|+pk+1}\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\theta+O(\delta^{kp}).$
@@
$p(k-|\alpha_{j}|)+1>-1,\quad 2p({2}|\alpha_{j}|-k)+p-4<0,$
where the first is required for integrability and the second is needed to obtain a meaningful estimate. The second is more restrictive. Now, if $\alpha_{j}=0$ we have
@@
$\int_{\Omega_{b_{j},\mathrm{I}}}\|\mathbf{X}_{n}(x,y)\|^{p}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y$ $\leq D_{\mathrm{I}}^{p}n^{-2pk+p-4}(\log n)^{2}+O(\delta^{kp}),$
@@
Then, again, rescaling and using polar coordinates and Lemma C.4,
$\int_{\Omega_{b_{j},\mathrm{II}}}\|\mathbf{X}_{n}(x,y)\|^{p}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y$ $\leq D_{\mathrm{II}}^{p}\left[N^{3p/2}n^{-4-2kp}\int_{0}^{N}r^{kp+1}\mathop{}\!\mathrm{d}r\right]+\mathrm{O}(\delta^{kp}),$
$=\frac{D_{\mathrm{II}}^{p}}{kp+2}N^{3p/2+kp+2}n^{-4-2kp}+\mathrm{O}(\delta^{kp}).$
@@
Suppose $p>2$ and $p(k-{2}|\alpha_{j}|)+2>0,p(k-{2}|\beta_{j}|)+2>0$ for all $j$. Suppose also that $\mathbf{V}_{n}(x,y)\in L^{\infty}(\mathbb{C})$ satisfies
@@
Then $\mathbf{V}_{n}(x,y)+\mathbf{I}$ is a solution of the $\overline{\partial}$ problem solved by $\mathbf{C}_{n}(x,y).$ Lastly, $\mathbf{V}_{n}$ is $\alpha$-Hölder continuous for some $0<\alpha<1$ and $\mathbf{V}_{n}|_{\Omega}$ is differentiable almost everywhere.
@@
We use [3, Theorem 4.3.10] which states that, in a distributional sense, the integral operator
$\mathcal{K}_{\Omega}u(x,y)=\frac{1}{\pi}\int_{\Omega}\frac{u(x^{\prime},y^{\prime})}{z^{\prime}-z}\mathop{}\!\mathrm{d}A(z^{\prime}),$
@@
$\overline{\partial}\mathbf{V}_{n}$ $=\mathbf{V}_{n}\mathbf{X}_{n}+\mathbf{X}_{n},$
$\overline{\partial}(\mathbf{V}_{n}+\mathbf{I})$ $=(\mathbf{V}_{n}+\mathbf{I})\mathbf{X}_{n}.$
Next, from [3, Theorem 4.3.13], we have that $\mathcal{K}_{\Omega}$ maps $L^{p}(\mathbb{C})$ into the space of $\alpha$-Hölder continuous functions for $\alpha=1-2/p$, $p>2$. And we again use the relation
$\mathbf{V}_{n}=\mathcal{K}_{\Omega}(\mathbf{V}_{n}\mathbf{X}_{n})+\mathcal{K}_{\Omega}\mathbf{X}_{n},$
to get the desired conclusion because $\mathbf{X}_{n}\in L^{p}(\Omega)$. The claim about $\mathbf{V}_{n}|_{\Omega}$ follows from the elliptic regularity theorem, see [23, Theorem 9.26], since $\mathbf{V}_{n}|_{\Omega}\in L^{\infty}(\Omega)$. ∎
We also see that for any $p>2>q,1/p+1/q=1$, and $\mathbf{U}\in L^{\infty}(\mathbb{C})$,
@@
Fix $k\geq 1$ and set $\eta={2}\max_{j}\max\{|\alpha_{j}|,|\beta_{j}|\}$. Suppose $2(\eta-k)-1<0$. Then for every $\gamma>0$, and $n$ sufficiently large, there is a unique $L^{\infty}(\mathbb{C})$ solution of (2.8) that satisfies
@@
By choosing $p>2$ in Lemma 2.3 sufficiently close to $2$, the theorem follows by Proposition 2.4. ∎
@@
Suppose $h_{j}\in C^{k,1}([a_{j},b_{j}])$ for each $j$. For $k\geq 1$ such that $2(\eta-k)<1$, $\eta={2}\max_{j}\max\{|\alpha_{j}|,|\beta_{j}|\}$ and any fixed $\gamma>0$, sufficiently small, set
$\mathbf{D}^{\natural}_{j}(z;c)$ $=\begin{bmatrix}d_{j,+}^{\sharp}(x,y;c)+1&0\\ 0&d_{j,-}^{\sharp}(x,y;c)+1\end{bmatrix}$,
$\mathbf{D}^{\natural}(x,y)$ $=\begin{cases}\mathbf{D}^{\natural}_{j}(z;a_{j})^{-1}\mathbf{D}^{\natural}_{j}(z;b_{j})^{-1},&z\in\Omega_{j,\pm},\\ \mathbf{I}&\text{otherwise},\end{cases}$
$\mathbf{F}^{\natural}(x,y)$ $=\begin{cases}\begin{bmatrix}1&0\\ \pm f_{j}^{\sharp}(x,y)/\omega_{j}(z)&1\end{bmatrix}&z\in\Omega_{j,\pm},\\ \mathbf{I}&\text{otherwise}.\end{cases}$
@@
All the estimates have been previously discussed and the remaining issue is the equality — that $\tilde{\mathbf{Y}}_{n}(z;\mu)=\mathbf{Y}_{n}(z;\mu)$. It can be shown that the jump conditions for $\tilde{\mathbf{R}}_{n}(z)$ satisfy the so-called product condition [36, Definition 2.55] and derivatives of the jump matrix are essentially bounded. Thus $\mathbf{O}_{n}\in H^{1}_{z}(\Gamma_{\delta})$, [36, Definition 2.48], an appropriate Sobolev space to ensure that $\tilde{\mathbf{R}}_{n}(z)$, as defined, is uniformly $\alpha$-Hölder continuous and takes $\alpha$-Hölder continuous boundary values for some $\alpha>0$. And $\tilde{\mathbf{Y}}_{n}(z;\mu)$ satisfies the jump conditions set for the FIK RH problem, in a continuous sense (away from $a_{j}$, $b_{j}$).
Next, we need to verify analyticity. But from classical elliptic regularity, [23, Theorem 9.26], we have that $\tilde{\mathbf{Y}}_{n}(\cdot;\mu)\in C^{\infty}(\Omega^{\prime})$ for every open set $\Omega^{\prime}$ that does not intersect any one of the contours used in the deformation, i.e. for
@@
Furthermore, $\overline{\partial}\tilde{\mathbf{Y}}_{n}=0$ in such a set $\Omega^{\prime}$, implying analyticity.
@@
So, if $n^{-1}\delta^{-3/2}\ll 1$, or $\delta=n^{-2/3+\gamma}$, and $2(\eta-k)+1/3<0$ we obtain a valid error term giving uniform asymptotics on scaled neighborhoods of the endpoints $a_{j},b_{j}$. Furthermore, by taking $k$ sufficiently large, $\gamma<2/3$ here can be chosen arbitrarily close to $2/3$.